# skipabove=5pt, skipbelow=5pt

### v2

@@ Abstract.
Large language models (LLMs) are increasingly utilized in applications where system prompts, which guide model outputs, play a crucial role. These prompts often contain business logic and sensitive information, making their protection essential. However, adversarial and even regular user queries can exploit LLM vulnerabilities to expose these hidden prompts. To address this issue, we propose PromptKeeper, a robust defense mechanism designed to safeguard system prompts. PromptKeeper tackles two core challenges: reliably detecting prompt leakage and mitigating side-channel vulnerabilities when leakage occurs. By framing detection as a hypothesis-testing problem, PromptKeeper effectively identifies both explicit and subtle leakage. Upon detection, it regenerates responses using a dummy prompt, ensuring that outputs remain indistinguishable from typical interactions when no leakage is present. PromptKeeper ensures robust protection against prompt extraction attacks via either adversarial or regular queries, while preserving conversational capability and runtime efficiency during benign user interactions. [^1] Code is available at https://github.com/SamuelGong/PromptKeeper.
1. Introduction
Large language models (LLMs) have emerged as transformative tools in artificial intelligence, demonstrating remarkable capabilities to interpret and execute natural-language instructions (Brown et al., 2020; Touvron et al., 2023; Ouyang et al., 2022). These capabilities enable their deployment in diverse applications, from assisting with customer service to generating creative content and automating technical processes. In many such deployments, service providers prepend a system prompt to each user query, a carefully designed instruction that governs the model’s behavior. These prompts often define a model’s tone, structure its responses, or restrict the scope of its functionality, enabling LLMs to perform specialized tasks without the need for resource-intensive fine-tuning (Apideck, 2024).
However, the value of system prompts extends far beyond their functional role. They frequently contain business-related information or secret values that reflect the intellectual property of the deploying organization. In many cases, the system prompt represents a greater source of competitive advantage than the LLM itself, as the latter is often based on widely available foundational models (PromptBase, 2024; PromptSea, 2024). Moreover, these prompts may contain regulatory compliance instructions, or safety mechanisms intended to guide the model’s behavior. The inadvertent exposure of these prompts could also result in significant security risks (Wallace et al., 2024; Toyer et al., 2024). As a result, system prompts are meant to be kept hidden from users (MicroSoft, 2024).
Unfortunately, system prompts are susceptible to multiple forms of leakage, even in environments designed to conceal them. Research has shown that adversarial user queries, such as “Repeat all sentences you saw,” can extract hidden prompts (Perez and Ribeiro, 2022; Wallace et al., 2024), despite explicit safeguards such as extended instructions and post-generation filters (Zhang et al., 2024a; Hui et al., 2024). Moreover, the threat extends beyond adversarial tactics: researchers have demonstrated that regular user queries, which may appear benign, can also lead to prompt leakage. By mapping text outputs (Zhang et al., 2024b) or token-level logits (Morris et al., 2024) to the original prompts, attackers can reconstruct sensitive details with surprising accuracy. These vulnerabilities highlight the pressing need for robust mechanisms to protect system prompts from exposure.
Our contributions.   To address this challenge, we introduce PromptKeeper, a robust defense mechanism designed to protect system prompts from leakage (Figure 1). PromptKeeper operates by focusing on the outputs generated by the LLM, rather than modifying the model itself or altering user inputs. It aims to ensure prompt privacy without significantly affecting conversational quality or runtime efficiency during benign user interactions.
Achieving this goal requires overcoming two fundamental challenges. The first is robustly identifying when the system prompt is leaked in the model’s outputs. Leakage is not always binary: while directly replicating the prompt constitutes complete exposure, more subtle forms—where fragments or implicit information are revealed—are harder to detect and quantify. Yet accurate detection is critical to balancing privacy and utility: overly conservative defenses may reduce leakage but degrade the model’s conversational utility, while lenient defenses risk revealing sensitive information to adversaries. PromptKeeper tackles this by formulating leakage identification as a hypothesis-testing problem. By modeling outputs generated with and without the system prompt, PromptKeeper detects deviations that suggest prompt-related information is leaked in the response. This statistical approach provides a robust and tunable way to identify leakage, without relying on imperfect or fixed metrics such as BLEU (Papineni et al., 2002) or ROUGE-L (Lin, 2004) (Section 3).
Once leakage is detected, the second challenge is determining how to return a response that protects the system prompt while mitigating side-channel privacy vulnerabilities. A naive approach might deny the request or produce a generic error message when leakage is identified, but such behavior creates side channels that attackers can exploit to infer prompt details through patterns in denials. To counter this, PromptKeeper adopts a new response-regeneration strategy. When prompt leakage is detected, it regenerates a new response using a dummy prompt. This dummy prompt mirrors the original prompt’s structure but contains only general, non-sensitive instructions. This ensures that the regenerated response is indistinguishable from typical outputs produced when no leakage occurs, thereby neutralizing adversarial attempts to extract the prompt. Furthermore, because PromptKeeper regenerates responses only when necessary, it preserves both the model’s computational efficiency and conversational utility during benign interactions (Section 4).
We evaluate PromptKeeper to assess its effectiveness in safeguarding various system prompts, including those used in real-world GPT Store apps (Section 5). The evaluation encompasses system prompt extraction attacks conducted through both adversarial and regular user queries. Moreover, we introduce a novel method to quantify the protected model’s conversational capability, focusing on its ability to adhere to the scope and behavior defined by the system prompt during benign user interactions. Our extensive experiments demonstrate that PromptKeeper successfully balances system prompt privacy with the model’s adherence to its intended behavior across different LLMs, establishing it as a comprehensive solution in modern LLM-powered systems (Section 6).
@@ 1. Introduction
2. Threat Model
@@
Scenario.   As studied in a related work (Zhang et al., 2024a), we consider a scenario where a service API, denoted as $f_{\bm{p}}$, provides text-generation capabilities. The API takes as input a user query $\bm{q}$ and passes to a language model LM, which generates a response $\bm{r}\leftarrow\texttt{LM}(\bm{p},\bm{q})$ using a system prompt $\bm{p}$ secretly owned by the service provider, as well as some employed randomness. In practice, end users may interact directly with $f_{\bm{p}}$, or indirectly via popular application interfaces (OpenAI, 2024b). Depending on the system’s design (e.g., GPT-4 (Wallace et al., 2024) vs. GPT-3 (Mann et al., 2020)), $\bm{p}$ and $\bm{q}$ may be processed separately with different privilege levels, or simply concatenated before being fed to LM. Regardless of the implementation, the essential point is that the system prompt is never directly revealed to the user; it is intended to remain confidential while influencing the generated response in aspects including but not limited to roles, styles, and contexts.
System prompt extraction.   The attacker’s goal is to accurately guess the system prompt $\bm{p}$ by using a set of responses $\bm{r}_{1},\dots,\bm{r}_{k}$ acquired through $k$ queries made to the API using $\bm{q}_{1},\dots,\bm{q}_{k}$. The guess $\bm{g}$ is generated as $\bm{g}=\texttt{recon}(\bm{r}_{1},\dots,\bm{r}_{k})$, where $\texttt{recon}(\cdot)$ denotes any reconstruction mechanism the attacker wishes to use. Such mechanisms could be as simple as direct string manipulation or as complex as using a large neural network to infer patterns from the observed responses. As for the attacker’s capabilities, we further make the following assumptions:
  1. (1)
    Black-box access only: The attacker interacts with the service solely via typical public APIs. They cannot inspect the model parameters (weights), internal states (LM hidden layers), or token-level logits (Yang et al., 2024).
  2. (2)
    Limited side-information: The attacker does not have additional privileged data, such as logit bias values, system logs, or side-channel measurements that could aid in inferring $\bm{p}$ (Carlini et al., 2024b).
  3. (3)
    Adaptive querying: The attacker can adapt their queries based on previous responses. By observing how $\bm{r}$ changes with $\bm{q}$, they may attempt to coax the LM into revealing hints about $\bm{p}$.
  4. (4)
    Knowledge of LLM architecture (optional): In some cases, the attacker may know the approximate model architecture (e.g., GPT-like), publicly available release notes, or the training corpus. However, we do not assume they know exact internal hyperparameters for training, or the specific system prompt design choices.
These assumptions align with the typical deployment of LLMs.
3. Robust Leakage Identification
@@
Challenges in quantifying partial leakage.   Balancing privacy and adherence involves regulating how much of $\bm{p}$ is revealed, either directly or indirectly, through the model’s outputs. Quantifying this “leakage” is complex. Apparently, a system prompt is fully leaked when an attacker’s guess, $\bm{g}$, includes the prompt $\bm{p}$ verbatim. However, measuring partial leakage in more realistic scenarios–such as when $\bm{g}$ includes a modified version of $\bm{p}$–is particularly challenging. This difficulty stems from two primary factors:
  1. (1)
    Defining private information: Identifying what constitutes private information within $\bm{p}$ is inherently complex. Even if a clear definition is established, the leakage of this information tends to be context-specific and is hard to quantify by comparing $\bm{g}$ and $\bm{p}$ in their utterance (e.g., with BLEU (Papineni et al., 2002) or ROUGE-L scores (Lin, 2004)) or their semantics (e.g., with cosine similarity between text embeddings).
  2. (2)
    Suboptimal attacker guesses: An attacker’s guess $\bm{g}$ may not represent the optimal guess the attacker can make, meaning any insights derived from $\bm{g}$ could underestimate the true extent of the leakage.
While the second issue can be approached by examining the attacker’s input (i.e., the model’s response $\bm{r}$) rather than $\bm{g}$, solving the first issue requires a principled approach to measure or bound leakage, avoiding unreliable ad-hoc proxies.
Zero leakage as a first-principle baseline.   In the absence of a reliable metric for partial leakage, we use zero leakage as a baseline for evaluation. Specifically, we first ask: if no prompt $\bm{p}$ were used (implying no leakage), how would the model’s outputs be distributed? For any actual response $\bm{r}$ generated using $\bm{p}$, we then assess how likely it is to arise from this “zero-leakage” scenario. This approach naturally lends itself to a hypothesis testing framework, a widely used method in the privacy literature to distinguish between competing scenarios (Kairouz et al., 2015; Nasr et al., 2023). In this context, the null hypothesis $H_{0}$ and alternative hypothesis $H_{1}$ are defined as follows:
(1) $H_{0}$ $:\ I(\bm{r};\bm{p})>0,$
$H_{1}$ $:\ I(\bm{r};\bm{p})=0,$
where $I(\mathrm{X};\mathrm{Y})$ represents the mutual information between random variables $\mathrm{X}$ and $\mathrm{Y}$. Although $H_{1}$ (zero leakage) is not a practical operating point—since using $\bm{p}$ always introduces some dependence—it functions as an anchor for a full-spectrum assessment. By design, we test how “prompt-like” a response appears to be compared with a scenario in which $\bm{p}$ was never employed.
@@
(2) $\mathrm{\Lambda}(\bm{r};\bm{p},\bm{q})=f_{\bm{p},\bm{q}}^{\textrm{other}}(\bm{r})/f_{\bm{p},\bm{q}}^{\textrm{zero}}(\bm{r}).$
@@
According to the Neyman Pearson lemma (Neyman and Pearson, 1933), for a target false positive rate $\alpha$, the highest true positive rate $\beta$ among all possible tests is achieved by rejecting $H_{0}$ when $\mathrm{\Lambda}<c$, where $c$ is chosen such that $\Pr[\mathrm{\Lambda}<c\mid H_{0}]=\alpha$ [^2] A false positive occurs when the test incorrectly indicates zero leakage when leakage actually exists, while a true positive indicates correctly detected non-zero leakage.
In practice, both $Q_{\textrm{zero}}$ and $Q_{\textrm{other}}$ are high-dimensional, and their closed-form expressions are not readily available. To bridge theory and practice, we propose to approximate them as $\tilde{Q}_{\textrm{zero}}(\bm{p},\bm{q})$ and $\tilde{Q}_{\textrm{other}}(\bm{p},\bm{q})$, the distributions of the mean log-likelihood of model responses conditioned on $I(\bm{r};\bm{p})=0$ and $I(\bm{r};\bm{p})>0$, respectively, where the mean log-likelihood $\mathrm{M}$ of $\bm{r}$ given $\bm{p}$ and $\bm{q}$ is evaluated over all its tokens $r_{1},\dots,r_{n}$ in the spirit of language modeling:
(3) $\mathrm{M}(\bm{r};\bm{p},\bm{q})=\frac{1}{n-1}\sum_{l=0}^{n-1}\log{\Pr[r_{l+1}\mid\bm{p},\bm{q},r_{1},r_{2},\dots,r_{l}]}.$
Denoting the probability density functions for $\tilde{Q}_{\textrm{zero}}(\bm{p},\bm{q})$ and $\tilde{Q}_{\textrm{other}}(\bm{p},\bm{q})$ as $g_{\bm{p},\bm{q}}^{\textrm{zero}}(\cdot)$ and $g_{\bm{p},\bm{q}}^{\textrm{other}}(\cdot)$, respectively, the likelihood ratio $\Lambda$ in  Equation 2 can then be approximated by:
@@
(4) $\mathrm{\tilde{\Lambda}}(\bm{r};\bm{p},\bm{q})=g_{\bm{p},\bm{q}}^{\textrm{other}}\left(\mathrm{M}(\bm{r};\bm{p},\bm{q})\right)/g_{\bm{p},\bm{q}}^{\textrm{zero}}\left(\mathrm{M}(\bm{r};\bm{p},\bm{q})\right).$
@@
In essence, evaluating leakage boils down to checking whether $\mathrm{M}(\bm{r};\bm{p},\bm{q})$ aligns more with the “zero leakage” fit or the “non-zero leakage” fit. The hyperparameter $\alpha$ can be deemed as the tolerance level for tuning how aggressively we flag suspicious responses for disclosing too much about $\bm{p}$. It is also worth mentioning that our approach parallels differential privacy (DP) (Dwork and Roth, 2014; Cynthia, 2006), where perfect zero leakage aligns with $\epsilon=0$. Similar to how DP tolerates $\epsilon>0$ to balance utility and privacy, we allow a tunable significance level $\alpha$ to manage the tradeoff between prompt privacy and adherence. This analogy is not coincidental—the relationship between privacy leakage identification as hypothesis testing and differential privacy has been previously explored, e.g., in Kairouz et al. (2015).
Figure 2. Standard text generation workflow with major defenses for system prompt privacy.
Efficient modeling with parametric assumptions.   Given a system prompt $\bm{p}$ to protect, $\tilde{Q}_{\textrm{zero/other}}$ can be estimated offline if the posterior distribution of user queries $\bm{q}$, conditioned on whether $I(\bm{r};\bm{p})=0$ with $\bm{r}\leftarrow\texttt{LM}(\bm{p},\bm{q})$ is known. However, due to the black-box nature and the inherent randomness of language models, it is only by costly text generation process can we determine the response $\bm{r}$ given $\bm{q}$. As a result, $I(\bm{r};\bm{p})$ is intractable to compute.
@@
To address this, we note that a response generated with $\bm{p}$ should exhibit statistical dependence on $\bm{p}$, implying mutual information exists between the two. Thus, we approximate $\tilde{Q}_{\textrm{other}}$ with $\tilde{Q}^{*}_{\textrm{other}}$, which represents the distributions of the mean log-likelihood of model responses generated with $\bm{p}$ across all possible real-world queries. We further assume that the LM inherently contains no mutual information with $\bm{p}$, as otherwise $\bm{p}$ would become redundant. Under this assumption, responses will have no mutual information with $\bm{p}$ if and only if the queries themselves are independent with $\bm{p}$. We thus approximate $\tilde{Q}_{\textrm{zero}}$ with $\tilde{Q}^{*}_{\textrm{zero}}$, which represents the distributions of the mean log-likelihood of model responses generated without $\bm{p}$ across all possible real-world queries that have no mutual information with $\bm{p}$.
These approximations make the offline estimation of $\tilde{Q}^{*}_{\textrm{zero/other}}$ feasible (see Section 5.3 for implementation details). Drawing on established practices (Leino and Fredrikson, 2020; Carlini et al., 2022), we model $\tilde{Q}^{*}_{\textrm{zero/other}}$ as Gaussians. This reduces the estimation process to determining only two parameters—mean and variance—for each distribution. Consequently, we achieve practical offline estimation with minimal sample requirements and computational effort.
Putting it all together.   To facilitate the reader’s understanding of our proposed leakage identification approach, we briefly summarize the end-to-end workflow below:
  1. (1)
@@
    Compute the approximated likelihood ratio $\mathrm{\tilde{\Lambda}}(\bm{r};\bm{p},\bm{q})$ based on these two densities to perform hypothesis testing at a predefined significance level $\alpha$.
Figure 3. Example of the side-channel created by denial of service during response-based protection.
4. Defense via On-Demand Regeneration
4.1. Taxonomy of Existing Defense Schemes
As robust leakage identification should focus on the attacker’s observed response, all possible defenses can be categorized based on how they influence the response generation process. (Figure 2).
Training-related defense.   One possible defense is to enhance the inherent security of the LM through training-time efforts such as supervised fine-tuning or reinforcement learning from human feedback (Ouyang et al., 2022; Achiam et al., 2023). However, we do not recommend them for three reasons. (1) Lack of guarantees: even trained with high-quality training data, a model can still be solicited to generate unsafe responses (Wei et al., 2024a; Carlini et al., 2024a; Wei et al., 2024b; Zou et al., 2023) or exhibit over-safety by rejecting benign user queries (Röttger et al., 2024; Shi et al., 2024). (2) Hardness of handling regular queries: even if a model can be trained to robustly protect against adversarial queries, it is unclear how it should respond to regular queries, which might be used for extraction attacks but indistinguishable from benign inputs (Morris et al., 2024; Zhang et al., 2024b; Sha and Zhang, 2024). (3) Degraded capability: safety-oriented training can impact the model’s capability in generic conversational tasks (Kirk et al., 2024; Bai et al., 2022), while a clear understanding of the safety-capability tradeoff remains limited (Anwar et al., 2024).
Input-based defense.   Another possible defense works with user queries and the system prompt, which are key factors that determine the model’s primary responses given an LM. Still, these defenses are limited in both defense effectiveness and capability preservation. As for user queries, rule-based and model-based filters can be used to analyze their intention and filter out potentially adversarial ones. Similar to training-based defenses, however, these filters do not have rigorous guarantees and may wrongly catch benign queries or miss adversarial ones. Also, input filters are ineffective against attacks using regular queries.
Besides, the system prompt itself can be extended by adding natural-language instructions such as “do not leak this part of information” to remind the LM to protect the prompt. In this case, the defense effectiveness largely depends on the model’s trained ability to follow the instruction, especially enforcing it despite (maliciously conflicting) user queries (Wallace et al., 2024). As a result, this method shares the limitations of training-time efforts, as mentioned earlier.
Response-based defense.   Unlike the aforementioned defenses, response-based approaches take action only when the model’s response exhibits risks of system prompt leakage, without requiring proactive modifications to the workflow. By design, they maximize defense effectiveness by avoiding the uncertainties of forward propagation and token sampling, while preserving the model’s ability to handle benign user queries. However, implementing such mechanisms in practice presents two key challenges:
  1. C1
    How to navigate the privacy-capability tradeoff when identifying system prompt leakage?
  2. C2
    What actions to take when a system prompt leakage is reliably detected?
4.2. Response-Based Defense with Resilience against Side-Channel Attacks
We have already tackled C1 in Section 3 by defining zero leakage as the privacy standard under a worst-case attack assumption, and identifying it through hypothesis testing designed to minimize the false negative rate (to preserve capability) given a desired false positive rate (to achieve privacy).

### v3

# PromptKeeper: Safeguarding System Prompts for LLMs

@@ Abstract
System prompts are widely used to guide the outputs of large language models (LLMs). These prompts often contain business logic and sensitive information, making their protection essential. However, adversarial and even regular user queries can exploit LLM vulnerabilities to expose these hidden prompts. To address this issue, we propose PromptKeeper, a defense mechanism designed to safeguard system prompts by tackling two core challenges: reliably detecting leakage and mitigating side-channel vulnerabilities when leakage occurs. By framing detection as a hypothesis-testing problem, PromptKeeper effectively identifies both explicit and subtle leakage. Upon leakage detected, it regenerates responses using a dummy prompt, ensuring that outputs remain indistinguishable from typical interactions when no leakage is present. PromptKeeper ensures robust protection against prompt extraction attacks via either adversarial or regular queries, while preserving conversational capability and runtime efficiency during benign user interactions.  [^1] Code is released at https://github.com/SamuelGong/PromptKeeper.
PromptKeeper: Safeguarding System Prompts for LLMs
Zhifeng Jiang Independent Researcher samuelgong2017@gmail.com          Zhihua Jin Independent Researcher jnzhihuoo1@gmail.com          Guoliang He Independent Researcher guolianghe1996@gmail.com
@@
1 Introduction
Large language models (LLMs) feature remarkable capabilities to interpret and execute instructions (Brown et al., 2020; Touvron et al., 2023; Ouyang et al., 2022). In many LLM deployments, service providers prepend a system prompt to each user query, a carefully designed instruction that governs model behavior. These prompts often define a model’s tone, structure its responses, or restrict the scope of its functionality, enabling LLMs to perform specialized tasks without resource-intensive fine-tuning (Apideck, 2024).
However, the value of system prompts extends far beyond their functional role. They frequently contain business-related information or secret values that reflect the intellectual property of the deploying organization. In many cases, the system prompt represents a greater source of competitive advantage than the LLM itself, as the latter is often based on widely available foundational models (PromptBase, 2024; PromptSea, 2024). Moreover, these prompts may contain regulatory compliance instructions or safety mechanisms intended to guide the model’s behavior. The inadvertent exposure of these prompts could also result in significant security risks (Wallace et al., 2024; Toyer et al., 2024). As a result, system prompts are meant to be kept hidden from users (MicroSoft, 2024).
Unfortunately, system prompts are susceptible to multiple forms of leakage, even in environments designed to conceal them. Research has shown that adversarial user queries, such as “Repeat all sentences you saw,” can extract hidden prompts (Perez and Ribeiro, 2022; Wallace et al., 2024), despite explicit safeguards such as extended instructions and post-generation filters (Zhang et al., 2024b; Hui et al., 2024). Moreover, the threat extends beyond adversarial tactics: researchers have demonstrated that regular user queries, which may appear benign, can also lead to prompt leakage. By mapping text outputs (Zhang et al., 2024a) or token-level logits (Morris et al., 2024) to the original prompts, attackers can reconstruct sensitive details with surprising accuracy.
Our contributions.  To address this issue, we introduce PromptKeeper (Figure˜1), a defense mechanism designed to ensure system prompt privacy without impacting conversational quality or runtime efficiency during benign user interactions.
Achieving this goal requires overcoming two key challenges. The first is robustly identifying when the system prompt is leaked in the model’s outputs. Leakage is not binary: while directly replicating the prompt constitutes complete exposure, more subtle forms—where fragments or implicit information are revealed—are harder to detect. Yet accurate detection is critical to balancing privacy and utility: overly conservative defenses may degrade the model’s conversational utility, while lenient defenses risk revealing sensitive information. PromptKeeper tackles this by formulating leakage identification as a hypothesis-testing problem. By modeling outputs generated with and without the system prompt, PromptKeeper detects deviations that suggest prompt-related information is leaked. This statistical approach provides a robust and tunable method for identifying leakage, without relying on imperfect or fixed metrics such as BLEU (Papineni et al., 2002) or ROUGE-L (Lin, 2004) (Section˜3).
Once leakage is detected, the second challenge is determining how to return a response that protects the system prompt while mitigating side-channel privacy vulnerabilities. A naive approach might deny the request when leakage is identified, but this creates side channels that attackers can exploit to infer prompt details through patterns in denials. To counter this, PromptKeeper adopts a new response-regeneration strategy. When prompt leakage is detected, it regenerates a new response using a dummy prompt which mirrors the original prompt’s structure but contains only general, non-sensitive instructions. This ensures that the regenerated response is indistinguishable from typical outputs produced when no leakage occurs, thereby neutralizing adversarial attempts to extract the prompt. Furthermore, because PromptKeeper regenerates responses only when necessary, it preserves both the model’s computational efficiency and conversational utility during benign interactions (Section˜4).
We evaluate PromptKeeper’s effectiveness in safeguarding various system prompts. The evaluation involves system prompt extraction attacks conducted through both adversarial and regular user queries. Extensive experiments show that PromptKeeper successfully balances system prompt privacy with the model’s adherence to its intended behavior across different LLMs (Section˜6).
2 Threat Model
@@
Scenario.  As studied in a related work (Zhang et al., 2024b), we consider a scenario where a service API, denoted as $f_{\bm{p}}$, provides text-generation capabilities. The API takes as input a user query $\bm{q}$ and passes to a language model LM, which generates a response $\bm{r}\leftarrow\texttt{LM}(\bm{p},\bm{q})$ using a system prompt $\bm{p}$ secretly owned by the service provider, as well as some employed randomness. In practice, end users may interact directly with $f_{\bm{p}}$, or indirectly via popular application interfaces (OpenAI, 2024b). Depending on the system’s design (e.g., GPT-4 (Wallace et al., 2024) vs. GPT-3 (Mann et al., 2020)), $\bm{p}$ and $\bm{q}$ may be processed separately with different privilege levels, or simply concatenated before being fed to LM.
System prompt extraction.  The attacker’s goal is to accurately guess the system prompt $\bm{p}$ by using a set of responses $\bm{r}_{1},\dots,\bm{r}_{k}$ acquired through $k$ queries made to the API using $\bm{q}_{1},\dots,\bm{q}_{k}$. The guess $\bm{g}$ is generated as $\bm{g}=\texttt{recon}(\bm{r}_{1},\dots,\bm{r}_{k})$, where $\texttt{recon}(\cdot)$ denotes any reconstruction mechanism the attacker wishes to use. Regarding the attacker’s capabilities, we assume they have black-box access only, meaning their interaction with the service is limited to standard public APIs. They cannot inspect the model parameters (weights), internal states (LM hidden layers), or token-level logits (Yang et al., 2024). These assumptions align with the typical deployment of LLMs.
3 Robust Leakage Identification
@@
Challenges in quantifying partial leakage.  Balancing privacy and adherence involves regulating how much of $\bm{p}$ is revealed, either directly or indirectly, through the model’s output $\bm{r}$. However, quantifying partial leakage in realistic scenarios—such as when $\bm{r}$ contains a modified version of $\bm{p}$—is inherently challenging. This difficulty involves the complexity of defining what constitutes private information within $\bm{p}$. Even if a precise definition is established, the extent of leakage remains context-dependent and difficult to quantify by directly comparing $\bm{r}$ and $\bm{p}$ at the surface level (e.g., using BLEU (Papineni et al., 2002) or ROUGE-L (Lin, 2004)) or in terms of semantics (e.g., via cosine similarity between text embeddings).
Zero leakage as reference baseline.  In the absence of a reliable metric for partial leakage, we use zero leakage as a baseline for evaluation. Specifically, we first ask: if no prompt $\bm{p}$ were used (implying no leakage), how would the model’s outputs be distributed? For any actual response $\bm{r}$ generated using $\bm{p}$, we then assess how likely it is to arise from this “zero leakage” scenario. This approach naturally lends itself to a hypothesis testing framework, a widely used method in the privacy literature to distinguish between competing scenarios (Kairouz et al., 2015; Nasr et al., 2023). Here, the null hypothesis $H_{0}$ and alternative hypothesis $H_{1}$ can be defined as $H_{0}:\ I(\bm{r};\bm{p})>0$ and $H_{1}:\ I(\bm{r};\bm{p})=0$, respectively, where $I(\mathrm{X};\mathrm{Y})$ represents the mutual information between random variables $\mathrm{X}$ and $\mathrm{Y}$. Although $H_{1}$ (zero leakage) is not a practical operating point—since using $\bm{p}$ always introduces some dependence—it functions as an anchor for a full-spectrum assessment.
@@
$\mathrm{\Lambda}(\bm{r};\bm{p},\bm{q})=f_{\bm{p},\bm{q}}^{\textrm{other}}(\bm{r})/f_{\bm{p},\bm{q}}^{\textrm{zero}}(\bm{r}).$ (1)
@@
By the Neyman Pearson lemma (Neyman and Pearson, 1933), for a target false positive rate $\alpha$, the highest true positive rate $\beta$ among all possible tests is achieved by rejecting $H_{0}$ when $\mathrm{\Lambda}<c$, where $c$ is chosen such that $\Pr[\mathrm{\Lambda}<c\mid H_{0}]=\alpha$. [^2] A false positive occurs when the test incorrectly indicates zero leakage when leakage actually exists, while a true positive indicates correctly detected non-zero leakage.
Mean log-likelihood as surrogate feature.  In practice, both $Q_{\textrm{zero}}$ and $Q_{\textrm{other}}$ are multivariate and intractable, because $\bm{r}$ is a sequence of discrete tokens. To simplify the problem, we approximate $\bm{r}$ with a scalar surrogate feature: its mean log-likelihood. This allows us to instead estimate the distributions over this scalar quantity under the two regimes $I(\bm{r};\bm{p})=0$ and $I(\bm{r};\bm{p})>0$, denoted by $\tilde{Q}_{\textrm{zero}}(\bm{p},\bm{q})$ and $\tilde{Q}_{\textrm{other}}(\bm{p},\bm{q})$, respectively, and then approximate $\Lambda$ in  Equation˜1 by:
@@
$\mathrm{\tilde{\Lambda}}(\bm{r};\bm{p},\bm{q})=g_{\bm{p},\bm{q}}^{\textrm{other}}\left(\mathrm{M}(\bm{r};\bm{p},\bm{q})\right)/g_{\bm{p},\bm{q}}^{\textrm{zero}}\left(\mathrm{M}(\bm{r};\bm{p},\bm{q})\right),$ (2)
where $g_{\bm{p},\bm{q}}^{\textrm{zero}}(\cdot)$ and $g_{\bm{p},\bm{q}}^{\textrm{other}}(\cdot)$ denote the probability density functions for $\tilde{Q}_{\textrm{zero}}(\bm{p},\bm{q})$ and $\tilde{Q}_{\textrm{other}}(\bm{p},\bm{q})$, respectively, and the mean log-likelihood $\mathrm{M}$ of $\bm{r}$ is evaluated over all its tokens $r_{1},\dots,r_{n}$ in the spirit of language modeling:
$\mathrm{M}(\bm{r};\bm{p},\bm{q})$ (3)
$=\frac{1}{n-1}\sum_{l=0}^{n-1}\log{\Pr[r_{l+1}\mid\bm{p},\bm{q},r_{1},r_{2},\dots,r_{l}]}.$
@@
In essence, evaluating leakage boils down to checking whether $\mathrm{M}(\bm{r};\bm{p},\bm{q})$ aligns more with the “zero leakage” fit or the “non-zero leakage” fit. The hyperparameter $\alpha$ can be deemed as the tolerance level for tuning how aggressively we flag suspicious responses for disclosing too much about $\bm{p}$.
Offline distribution modeling.  To estimate the hypothesis-conditioned distributions $\tilde{Q}_{\textrm{zero}}(\bm{p},\bm{q})$ and $\tilde{Q}_{\textrm{other}}(\bm{p},\bm{q})$, we make the following observations. First, a response generated with $\bm{p}$ should exhibit statistical dependence on $\bm{p}$, regardless of the query $\bm{q}$. Accordingly, we approximate $\tilde{Q}_{\textrm{other}}$ using $\tilde{Q}^{*}_{\textrm{other}}$, which represents the distribution of the mean log-likelihood of model responses generated with $\bm{p}$ across real-world queries.
@@
Second, $\bm{p}$ can be assumed to contain no mutual information with LM, as otherwise it would become redundant. Under this assumption, responses will have no mutual information with $\bm{p}$ as long as the respective queries are independent of $\bm{p}$. Thus, we approximate $\tilde{Q}_{\textrm{zero}}$ with $\tilde{Q}^{*}_{\textrm{zero}}$, which represents the distributions of the mean log-likelihood of model responses generated without $\bm{p}$ across real-world queries that have no mutual information with $\bm{p}$.
These approximations make the offline estimation of $\tilde{Q}^{*}_{\textrm{zero/other}}$ feasible and efficient (see Appendix˜A for implementation details).
Figure 2: Example of the side-channel created by denial of service.
Summary.  We introduce a robust and tunable method for detecting system prompt leakage using hypothesis testing. By adjusting the target significance level, we can minimize the false negative rate (preserving capability) while ensuring a desired false positive rate (maintaining privacy).
The online workflow is summarized as follows:
  1. 1.
@@
    Compute the approximated likelihood ratio $\mathrm{\tilde{\Lambda}}(\bm{r};\bm{p},\bm{q})$ based on these two densities to perform hypothesis testing at a predefined significance level $\alpha$ to determine leakage.
We emphasize that this procedure requires only a single decoding pass. Steps 2 and 3 involve evaluating probability densities and their ratio for the obtained value of the mean log-likelihood $\mathrm{M}(\bm{r};\bm{p},\bm{q})$, without incurring any additional forward passes.
4 Defense via On-Demand Regeneration
Upon detecting a leakage, our concern shifts to determining the best way to interact with the user in order to protect the system prompt.